What is the capacity of a gas tank (in gal) if it contains 125 lb of gas. The density of the gas is 0.0071 g/ml.(1 lb=454.3 g, 1 gal = 3.785 L )
Answer #1

Mass of gas = 125 lb

                      = 125 * 454.3 g

                     = 56787.5 g

Density of gas = 0.0071 g/mL

Volume of gas = mass / density

                          = 56787.5 g / 0.0071 g/mL

                          = 7998239.437 mL

                          = 7998239.437 * 10-3 L                         [1mL = 10-3 L]

                          = 7998239.437 * 10-3/ 3.785 gal

                          = 2113.14 gal

Volume of gas is equal to the capacity of tank. So capacity of tank is 2113. 14 gal